# Revisiting Non-Acyclic GFlowNets in Discrete Environments

### v2

@@ Appendix A Proofs
$11=\frac{\sum_{s^{\prime}\in\operatorname{out}(s_{t})}C(s^{\prime})\mathcal{F}(s_{t}\to s^{\prime})}{C(s_{t})\mathcal{F}(s_{t})}>\frac{\sum_{s^{\prime}\in\operatorname{out}(s_{t})}C(s_{t})\mathcal{F}(s_{t}\to s^{\prime})}{C(s_{t})\mathcal{F}(s_{t})}=\frac{\sum_{s^{\prime}\in\operatorname{out}(s_{t})}\mathcal{F}(s_{t}\to s^{\prime})}{\mathcal{F}(s_{t})}=$ 1 .
Thus, in this case, there is also a contradiction. Therefore $C(s)=1$ is a unique solution, meaning that ^$\hat{\mathcal{F}}(s)=\mathcal{F}(s)$. Finally for any $s\to s^{\prime}\in\mathcal{E}$
^$\hat{\mathcal{F}}(s\to s^{\prime})=C(s^{\prime})\mathcal{F}(s\to s^{\prime})=\mathcal{F}(s\to s^{\prime}).$
@@
Uniqueness. Suppose that such a forward policy $\mathcal{P}_{\mathrm{F}}$ exists, then probability distributions over $\mathcal{T}$ induced by $\mathcal{P}_{\mathrm{B}}$ and $\mathcal{P}_{\mathrm{F}}$ coincide. Then
$\begin{split}\mathcal{F}(s\to s^{\prime})&=\mathcal{F}(s_{f})\mathbb{E}_{n_{\tau}}\mathopen{}\mathclose{{}\left[\mathbb{E}_{\tau}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}}\mathbb{I}\{s_{t}=s,s_{t+1}=s^{\prime}\}\mid n_{\tau}}\right]}\right]\\ &=\mathcal{F}(s_{f})\mathbb{E}_{n_{\tau}}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}}\mathbb{E}_{\tau}\mathopen{}\mathclose{{}\left[\mathbb{I}\{s_{t}=s,s_{t+1}=s^{\prime}\}\mid n_{\tau}}\right]}\right]\\ &=\mathcal{F}(s_{f})\mathbb{E}_{n_{\tau}}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}}\mathbb{P}\mathopen{}\mathclose{{}\left(s_{t}=s,s_{t+1}=s^{\prime}\mid n_{\tau}}\right)}\right]\\ &=\mathcal{F}(s_{f})\mathbb{E}_{n_{\tau}}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}}\mathbb{P}\mathopen{}\mathclose{{}\left(s_{t}=s\mid n_{\tau}}\right)\mathbb{P}\mathopen{}\mathclose{{}\left(s_{t+1}=s^{\prime}\mid s_{t}=s,n_{\tau}}\right)}\right]\\ &=\mathcal{F}(s_{f})\mathbb{E}_{n_{\tau}}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}}\mathbb{P}\mathopen{}\mathclose{{}\left(s_{t}=s\mid n_{\tau}}\right)}\right]\mathcal{P}_{\mathrm{F}}(s^{\prime}\mid s)=\mathcal{F}(s)\mathcal{P}_{\mathrm{F}}(s^{\prime}\mid s).\\ \end{split}$
Then we have $\mathcal{F}(s^{\prime})\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})=\mathcal{F}(s)\mathcal{P}_{\mathrm{F}}(s^{\prime}\mid s)$ and $\mathcal{P}_{\mathrm{F}}(s^{\prime}\mid s)=\mathcal{F}(s^{\prime})\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})/\mathcal{F}(s)$, thus we finish the proof of uniqueness presented above.
@@
where the first equation is due to Proposition 3.6. Next, note that
$\mathcal{F}(s_{0})=\mathcal{F}(s_{f})\;\mathbb{E}_{\tau}\mathopen{}\mathclose{{}\left[\sum_{t=0}^{n_{\tau}+1}\mathbb{I}\{s_{t}=s_{0}\}}\right]=\mathcal{F}(s_{f})\cdot 1=\mathcal{F}(s_{f}).$
Then
@@
Thus the existence is proven. Finally, the detailed balance conditions follow from the proof of uniqueness.
@@
Consider an edge function ¯$\bar{\mathcal{F}}(s\to s^{\prime})=\mathcal{F}(s)\mathcal{P}_{\mathrm{F}}(s^{\prime}\mid s)$. It is positive since $\mathcal{F}(s)>0$ and $\mathcal{P}_{\mathrm{F}}(s^{\prime}\mid s)>0$ by the statement of the proposition. Since $\mathcal{P}_{\mathrm{F}}(\cdot\mid s)$ is a valid probability distribution over $\operatorname{out}(s)$, we have
@@
Assume that (i) a graph $\mathcal{G}$ satisfies 3.1 and (ii) for any $s\in\mathcal{S},s^{\prime}\in\mathcal{A}_{s}$ it holds $r(s,s^{\prime})\leq 0$ and $r(s,s^{\prime})=0$ if and only if $|\operatorname{in}(s^{\prime})|=1$. Also, assume that for any optimal policy $\pi^{\star}$ it holds $\mathbb{E}_{\pi^{\star}}[n_{\tau}]<+\infty$.
@@ B.1 Training Policy and Flow Weighting
This implies that on-policy training tries to minimize the sum of squared state flows rather than the sum of state flows. This happens due to the fact that the trajectory distribution that is used to collect data for training (induced by $\mathcal{P}_{\mathrm{F}}$ in this case) visits certain states more often than others, thus a weight is given to the flow in each state equal to the expected number of visits. However, if $\mathcal{P}_{\mathrm{F}}(s\mid s_{0})$ is fixed to be uniform over $\mathcal{S}\setminus\{s_{0},s_{f}\}$ (see Section 4 and Appendix B.3), this issue can be circumvented by applying flow regularizer only in the first state of each sampled trajectory. Then, equal weight will be given to $\mathcal{F}_{\theta}(s)$ in each state in the expected loss, thus we will be minimizing the sum of state flows. However, in our experiments we noticed that this does not significantly influence the results, so we leave exploring this phenomenon as further research direction.
@@ B.2 Loss Scaling and Stability
In this section, we provide a more detailed explanation of our scaling hypothesis (see Section 4). Let us consider a GFlowNet that learns $\mathcal{F}$, $\mathcal{P}_{\mathrm{F}}$ and $\mathcal{P}_{\mathrm{B}}$. Since these quantities are predicted by a neural network, a standard way is to make it predict logits for the forward policy, logits for the backward policy, and the logarithm of the state flow. Flow functions are always positive, thus predicting them in log scale is a natural approach (Bengio et al., 2021, 2023). Then, for any transition $s\to s^{\prime}$, define two quantities:
@@
However, for both losses, one can either replace $\Delta_{\log\mathcal{F}}$ with $\Delta_{\mathcal{F}}$ or the other way around. For visualization, let us fix the predicted log backward flow $\mathcal{F}_{B}$ to be, e.g., $1$, and plot the losses with respect to the varying value of the predicted log forward flow $\mathcal{F}_{F}$. The plots are presented in Figure 3. One can note that as argument $\log\mathcal{F}_{F}$ decreases, both losses in $\Delta_{\mathcal{F}}$ scale quickly plateau, thus their derivative goes to zero. From the optimization perspective, this means that when the predicted log flow needs to be increased, the gradient step will be very small since the derivative of the loss is almost zero. On the other hand, when the predicted log flow needs to be decreased, the gradient step will be larger since losses have much higher derivatives in the corresponding regions. In combination with Proposition 3.12, this gives a possible explanation to the stability of $\Delta_{\mathcal{F}}$ scale losses: they are biased towards underestimation of the flows, and, as a result, biased towards solutions with smaller expected trajectory length. We note that the same reasoning can be applied to the stable flow matching loss proposed in (Brunswic et al., 2024) since it also operates with differences between flows in $\Delta_{\mathcal{F}}$ scale.
@@
Figure 3: Plots for DB and SDB losses in $\Delta\mathcal{F}$ and $\Delta\log\mathcal{F}$ scales with fixed predicted log backward flow $=1$ and varying predicted log forward flow. More specifically, green curve is $y=(x-1)^{2}$, red curve is $y=\mathopen{}\mathclose{{}\left(e^{x}-e^{1}}\right)^{2}$, brown curve is $y=\log\mathopen{}\mathclose{{}\left(1+(x-1)^{2}}\right)\cdot(1+0.001e^{x})$, blue curve is $y=\log\mathopen{}\mathclose{{}\left(1+\mathopen{}\mathclose{{}\left(e^{x}-e^{1}}\right)^{2}}\right)\cdot(1+0.001e^{x})$
@@ B.4 Solving Small Environments Exactly
Suppose we have a fixed backward policy $\mathcal{P}_{\mathrm{B}}$ and a final flow $\mathcal{F}(s_{f})$. Then, induced flows $\mathcal{F}$ and the corresponding forward policy $\mathcal{P}_{\mathrm{F}}$ can be obtained exactly for small environments. Consider the following system of linear equations with respect to ^$\hat{\mathcal{F}}(s)$ that arises from Proposition 3.6:
{^$\mathopen{}\mathclose{{}\left\{\begin{array}[]{l}\hat{\mathcal{F}}(s)=\sum_{s^{\prime}\in\operatorname{out}(s)}\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})\hat{\mathcal{F}}(s^{\prime}),\;\forall s\in\mathcal{S}\setminus\{s_{f}\},\\ \hat{\mathcal{F}}(s_{f})=\mathcal{F}(s_{f}).\end{array}}\right.$ (17)
The system has $|\mathcal{S}|$ variables and $|\mathcal{S}|$ equations. ^$\hat{\mathcal{F}}(s)=\mathcal{F}(s)$ is a solution, where $\mathcal{F}(s)$ are state flows induced by $\mathcal{P}_{\mathrm{B}}$ and $\mathcal{F}(s_{f})$, and the uniqueness of the solution follows from Proposition 3.7. Thus, by solving the system, one can exactly find induced state flows. Then, by Proposition 3.6 and Proposition 3.8, edge flows and $\mathcal{P}_{\mathrm{F}}$ can also be exactly expressed as
@@
Interestingly, the system (17) can also be explained from Markov Chain perspective. Let us take the graph $\mathcal{G}$ with reversed edges, add a loop from $s_{0}$ to itself, and use $\mathcal{P}_{\mathrm{B}}$ to define a Markov Chain with the following transition matrix: $P(s_{0}\mid s_{0})=1$, $P(s\mid s^{\prime})=\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})$ if there is an edge $s\to s^{\prime}$, and $P(s\mid s^{\prime})=0$ otherwise. It will be an absorbing Markov Chain, with an only absorbing state $s_{0}$ since it is reachable from any other state by Assumption 3.1. The transition matrix can be written in the following way:
@@
where $Q$ is a $|\mathcal{S}|-1$ by $|\mathcal{S}|-1$ matrix and $R$ is a $|\mathcal{S}|-1$ by $1$ matrix. Its fundamental matrix $N$, i.e., such matrix that $N_{s,s^{\prime}}$ is equal to the expected number of visits to a non-absorbing state $s^{\prime}$ before being absorbed when starting from a non-absorbing state $s$, can be obtained as:
@@
where $I-Q$ is always invertible (Kemeny & Snell, 1969, Theorem 3.2.1). One can note that normalized flows $\mathcal{F}(s)/\mathcal{F}(s_{f})$ coincide with the expected number of visits to $s$ when starting from $s_{f}$, thus coincide with the row of matrix $N$ corresponding to $s_{f}$. Finally, notice that $(I-Q)$ coincides with the transposed matrix of the truncated system (17) (with the exception of the variable and the equation corresponding to $s_{0}$), thus such system has a unique solution $\mathcal{F}(s_{f})(I-Q)^{-T}e_{s_{f}}=\mathcal{F}(s_{f})N^{T}e_{s_{f}}$, where $e_{s_{f}}$ is a vector of size $|\mathcal{S}|-1$ that has $1$ on the position corresponding to $s_{f}$ and 00 on all others. The variable corresponding to $s_{0}$ should be handled separately, but it is easy to see ^$\hat{\mathcal{F}}(s_{0})=\sum_{s^{\prime}\in\operatorname{out}(s)}\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime}){\mathcal{F}}(s^{\prime})=\mathcal{F}(s_{0})$.
@@ C.3.1 Reward Distribution Properties
$\begin{split}\mathopen{}\mathclose{{}\left|S_{1}\cup\cdots\cup S_{n}}\right|&=\sum_{i}\mathopen{}\mathclose{{}\left|S_{i}}\right|-\sum_{i<j}\mathopen{}\mathclose{{}\left|S_{i}\cap S_{j}}\right|+\sum_{i<j<k}\mathopen{}\mathclose{{}\left|S_{i}\cap S_{j}\cap S_{k}}\right|+\cdots+(-1)^{n+1}\mathopen{}\mathclose{{}\left|S_{1}\cap\cdots\cap S_{n}}\right|\\ &=\binom{n}{1}(n-1)!-\binom{n}{2}(n-2)!+\binom{n}{3}(n-3)!-\cdots+(-1)^{n+1}\binom{n}{n}0!\\ &=\sum_{i=1}^{n}(-1)^{i+1}\binom{n}{i}(n-i)!=n!\sum_{i=1}^{n}\frac{(-1)^{i+1}}{i!}.\\ \end{split}$

### v3

@@ Appendix A Proofs
$1=\frac{\sum_{s^{\prime}\in\operatorname{out}(s_{t})}C(s^{\prime})\mathcal{F}(s_{t}\to s^{\prime})}{C(s_{t})\mathcal{F}(s_{t})}>\frac{\sum_{s^{\prime}\in\operatorname{out}(s_{t})}C(s_{t})\mathcal{F}(s_{t}\to s^{\prime})}{C(s_{t})\mathcal{F}(s_{t})}=\frac{\sum_{s^{\prime}\in\operatorname{out}(s_{t})}\mathcal{F}(s_{t}\to s^{\prime})}{\mathcal{F}(s_{t})}=1.$
Thus, in this case, there is also a contradiction. Therefore $C(s)=1$ is a unique solution, meaning that $\hat{\mathcal{F}}(s)=\mathcal{F}(s)$. Finally for any $s\to s^{\prime}\in\mathcal{E}$
$\hat{\mathcal{F}}(s\to s^{\prime})=C(s^{\prime})\mathcal{F}(s\to s^{\prime})=\mathcal{F}(s\to s^{\prime}).$
@@
Uniqueness. Suppose that such a forward policy $\mathcal{P}_{\mathrm{F}}$ exists, then the probability distributions over $\mathcal{T}$ induced by $\mathcal{P}_{\mathrm{B}}$ and $\mathcal{P}_{\mathrm{F}}$ coincide. Moreover, since for any edge $s\to s^{\prime}\in\mathcal{E}$, we have $\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})>0$ and the probability distribution over trajectories coincide, we get that $\mathcal{P}_{\mathrm{F}}(s^{\prime}\mid s)>0$. Similarly to the proofs of Lemma 3.4 and Proposition 3.6, define a Markov chain $\{X_{n}\}_{n=0}^{\infty}$ corresponding to a random walk on $\mathcal{G}$ in the forward direction with transition probabilities given by $\mathcal{P}_{\mathrm{F}}$. Formally, $X_{0}=s_{0}$ a.s., $\mathbb{P}[X_{t}=s^{\prime}\mid X_{t-1}=s]=\mathcal{P}_{\mathrm{F}}(s^{\prime}\mid s)$, and $\mathbb{P}[X_{t}=s_{f}\mid X_{t-1}=s_{f}]=1$.
Similarly to the proof of Proposition 3.6, we have
[…]
In (a) we used the finiteness of the expected value to apply Fubini’s theorem, and in (b) we used the Markov property of $\{X_{n}\}_{n=0}^{\infty}$. Then, combined with Proposition 3.6, we have $\mathcal{F}(s^{\prime})\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})=\mathcal{F}(s)\mathcal{P}_{\mathrm{F}}(s^{\prime}\mid s)$. Thus, $\mathcal{P}_{\mathrm{F}}(s^{\prime}\mid s)=\mathcal{F}(s^{\prime})\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})/\mathcal{F}(s)$, concluding the proof of uniqueness.
@@
where the first equation is due to Proposition 3.6. By Proposition 3.6 we also have $\mathcal{F}(s_{0})=\mathcal{F}(s_{f})$. Then
@@
Thus the existence is proven. Finally, the detailed balance conditions follow from the proof of uniqueness presented above.
@@
Consider an edge function $\bar{\mathcal{F}}(s\to s^{\prime})=\mathcal{F}(s)\mathcal{P}_{\mathrm{F}}(s^{\prime}\mid s)$. It is positive since $\mathcal{F}(s)>0$ and $\mathcal{P}_{\mathrm{F}}(s^{\prime}\mid s)>0$ by the statement of the proposition. Since $\mathcal{P}_{\mathrm{F}}(\cdot\mid s)$ is a valid probability distribution over $\operatorname{out}(s)$, we have
@@
Assume that (i) a graph $\mathcal{G}$ satisfies Assumption 3.1 and (ii) for any $s\in\mathcal{S},s^{\prime}\in\mathcal{A}_{s}$ it holds $r(s,s^{\prime})\leq 0$ and $r(s,s^{\prime})=0$ if and only if $|\operatorname{in}(s^{\prime})|=1$. Also, assume that for any optimal policy $\pi^{\star}$ it holds $\mathbb{E}_{\pi^{\star}}[n_{\tau}]<+\infty$.
@@ B.1 Training Policy and Flow Weighting
This implies that on-policy training tries to minimize the sum of squared state flows rather than the sum of state flows. This happens due to the fact that the trajectory distribution that is used to collect data for training (induced by $\mathcal{P}_{\mathrm{F}}$ in this case) visits certain states more often than others, thus a weight is given to the flow in each state equal to the expected number of visits. However, if $\mathcal{P}_{\mathrm{F}}(s\mid s_{0})$ is fixed to be uniform over $\mathcal{S}\setminus\{s_{0},s_{f}\}$ (see Section 4 and Appendix B.3), this issue can be circumvented by applying the flow regularizer only in the first state of each sampled trajectory. Then, equal weight will be given to $\mathcal{F}_{\theta}(s)$ in each state in the expected loss, thus we will be minimizing the sum of state flows. However, in our experiments we noticed that this does not significantly influence the results, so we leave exploring this phenomenon as a further research direction.
@@ B.2 Loss Scaling and Stability
In this section, we provide a more detailed explanation of our scaling hypothesis (see Section 4). Let us consider a GFlowNet that learns $\mathcal{F}$, $\mathcal{P}_{\mathrm{F}}$ and $\mathcal{P}_{\mathrm{B}}$. Since these quantities are predicted by a neural network, a standard way is to make it predict logits for the forward policy, logits for the backward policy, and the logarithm of the state flow. Flow functions are always positive, thus, predicting them in log scale is a natural approach (Bengio et al., 2021, 2023). Then, for any transition $s\to s^{\prime}$, define two quantities:
@@
However, for both losses, one can either replace $\Delta_{\log\mathcal{F}}$ with $\Delta_{\mathcal{F}}$ or the other way around. For visualization, let us fix the predicted log backward flow $\mathcal{F}_{B}$ to be, e.g., $1$, and plot the losses with respect to the varying value of the predicted log forward flow $\mathcal{F}_{F}$. The plots are presented in Figure 3. One can note that as argument $\log\mathcal{F}_{F}$ decreases, both losses in $\Delta_{\mathcal{F}}$ scale quickly plateau, thus their derivative goes to zero. From the optimization perspective, this means that when the predicted log flow needs to be increased, the gradient step will be very small since the derivative of the loss is almost zero. On the other hand, when the predicted log flow needs to be decreased, the gradient step will be larger since losses have much higher derivatives in the corresponding regions. In combination with Proposition 3.12, this gives a possible explanation to the stability of $\Delta_{\mathcal{F}}$ scale losses: they are biased towards underestimation of the flows and, as a result, biased towards solutions with smaller expected trajectory length. We note that the same reasoning can be applied to the stable flow matching loss proposed in (Brunswic et al., 2024) since it also operates with differences between flows in $\Delta_{\mathcal{F}}$ scale.
@@
Figure 3: Plots for DB and SDB losses in $\Delta\mathcal{F}$ and $\Delta\log\mathcal{F}$ scales with fixed predicted log backward flow =1=1 and varying predicted log forward flow. More specifically, green curve is $y=(x-1)^{2}$, red curve is $y=\mathopen{}\mathclose{{\left(e^{x}-e^{1}}}\right)^{2}$, brown curve is $y=\log\mathopen{}\mathclose{{\left(1+(x-1)^{2}}}\right)\cdot(1+0.001e^{x})$, blue curve is $y=\log\mathopen{}\mathclose{{\left(1+\mathopen{}\mathclose{{\left(e^{x}-e^{1}}}\right)^{2}}}\right)\cdot(1+0.001e^{x})$
@@ B.4 Solving Small Environments Exactly
Suppose we have a fixed backward policy $\mathcal{P}_{\mathrm{B}}$ and a final flow $\mathcal{F}(s_{f})$. Then, induced flows $\mathcal{F}$ and the corresponding forward policy $\mathcal{P}_{\mathrm{F}}$ can be obtained exactly for small environments. Consider the following system of linear equations with respect to $\hat{\mathcal{F}}(s)$ that arises from Proposition 3.6:
$\mathopen{}\mathclose{{\left\{\begin{array}[]{l}\hat{\mathcal{F}}(s)=\sum_{s^{\prime}\in\operatorname{out}(s)}\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})\hat{\mathcal{F}}(s^{\prime}),\;\forall s\in\mathcal{S}\setminus\{s_{f}\},\\ \hat{\mathcal{F}}(s_{f})=\mathcal{F}(s_{f}).\end{array}}}\right$. (17)
The system has $|\mathcal{S}|$ variables and $|\mathcal{S}|$ equations. $\hat{\mathcal{F}}(s)=\mathcal{F}(s)$ is a solution, where $\mathcal{F}(s)$ are state flows induced by $\mathcal{P}_{\mathrm{B}}$ and $\mathcal{F}(s_{f})$, and the uniqueness of the solution follows from Proposition 3.7. Thus, by solving the system, one can exactly find induced state flows. Then, by Proposition 3.6 and Proposition 3.8, edge flows and $\mathcal{P}_{\mathrm{F}}$ can also be exactly expressed as
@@
Interestingly, the system (17) can also be explained from the Markov Chain perspective. Let us take the graph $\mathcal{G}$ with reversed edges, add a loop from $s_{0}$ to itself, and use $\mathcal{P}_{\mathrm{B}}$ to define a Markov Chain with the following transition matrix: $P(s_{0}\mid s_{0})=1$, $P(s\mid s^{\prime})=\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})$ if there is an edge $s\to s^{\prime}$, and $P(s\mid s^{\prime})=0$ otherwise. It will be an absorbing Markov Chain, with an only absorbing state $s_{0}$ since it is reachable from any other state by Assumption 3.1. The transition matrix can be written in the following way:
@@
where $Q$ is a $|\mathcal{S}|-1$ by $|\mathcal{S}|-1$ matrix and $R$ is a $|\mathcal{S}|-1$ by $1$ matrix. Its fundamental matrix $N$, i.e., such a matrix that $N_{s,s^{\prime}}$ is equal to the expected number of visits to a non-absorbing state $s^{\prime}$ before being absorbed when starting from a non-absorbing state $s$, can be obtained as:
@@
where IQI-Q is always invertible (Kemeny & Snell, 1969, Theorem 3.2.1). One can note that normalized flows $\mathcal{F}(s)/\mathcal{F}(s_{f})$ coincide with the expected number of visits to $s$ when starting from $s_{f}$, thus coincide with the row of matrix $N$ corresponding to $s_{f}$. Finally, notice that (IQ)(I-Q) coincides with the transposed matrix of the truncated system (17) (with the exception of the variable and the equation corresponding to $s_{0}$), thus, such system has a unique solution $\mathcal{F}(s_{f})(I-Q)^{-T}e_{s_{f}}=\mathcal{F}(s_{f})N^{T}e_{s_{f}}$, where $e_{s_{f}}$ is a vector of size $|\mathcal{S}|-1$ that has $1$ on the position corresponding to $s_{f}$ and 0 on all others. The variable corresponding to $s_{0}$ should be handled separately, but it is easy to see $\hat{\mathcal{F}}(s_{0})=\sum_{s^{\prime}\in\operatorname{out}(s)}\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime}){\mathcal{F}}(s^{\prime})=\mathcal{F}(s_{0})$.
@@ C.3.1 Reward Distribution Properties
[…]